OpenStudy (anonymous):

A sporting goods store has added a section with trail foods and snacks. Raisins cost $5.00 per pound and peanuts cost $3.80 per pound. Raisins and peanuts are going to be sold in a mixture of 30 pounds which will cost $4.05 per pound? How many pounds of raisins will be in the mix?

OpenStudy (radar):

Let r = pounds of raisins. then 30-r = pounds of peanuts. Now to work in the costs. 5r + (30-r)3.80 = 30(4.05) 5r+114-3.80r=121.5 solve for r 1.2r=121.5-114 1.2r=7.5 r=6.25 lbs peanuts =23.75 lbs.

OpenStudy (radar):

Probably should verify by checking. 6.25($5.00) +23.75($3.80) = 30($4.05) 31.25 +90.25=121.5 121.5=121.5 Answer is correct.